# Nonradial oscillations of stratified neutron stars with solid crusts: Mode characterization and tidal resonances in coalescing binaries

### v1

@@ I Introduction
As the densest stable form of matter known in the Universe, neutron stars (NSs) are expected to have complex interiors dominated by strong interactions inaccessible through terrestrial experiments. Despite extensive experimental and theoretical efforts in probing the nuclear equation of state (EOS), our current knowledge about the EOS at the densities most relevant for the determining the internal structure of NSs remains poor [1, 2, 3].
@@
Compact binary mergers involving NSs provide multiple pathways to probe the dense matter, from subtle finite-size effects in the late inspiral [4, 5, 6, 7, 6, 8, 9, 10, 11, 12, 13, 14] to the dramatic dynamics of merger remnants [15, 16, 17, 18, 19, 20], tidal disruption [21, 22, 23], and mass ejection [24, 25, 26]. The landmark detection of gravitational waves (GWs) from the binary NS inspiral GW170817 by the LIGO-Virgo-KAGRA Collaboration [7], along with its postmerger electromagnetic (EM) counterparts [27], marked the beginning of GW multimessenger astronomy, and has helped shed a new light on the EOS problem. However, the constraints on the EOS set by GW170817 are limited by the insufficient sensitivity of current GW detectors at the frequency band of the late-inspiral and merger, where matter effects on the GW waveform become most salient.
@@
An upgrade to the current ground-based detectors is pending for the coming years, and the next observing run of the International Gravitational-Wave Observatory Network, tentatively planned for 2028, will provide an opportunity to detect GWs generated during the late-inspiral and merger of GW170817-like events. A central question to be addressed is then: To what extent can observations of binary mergers truly reveal the composition and state of matter inside NSs? Because a NS’s characteristic oscillation frequencies are determined by its internal structure, identifying quasiperiodic oscillations (QPOs) in some observations offers a way to probe these extreme environments. This article focuses on a subset of the stellar spectrum that may imprint observable signatures in gravitational waveforms or EM flares through tidal interactions.
@@
Several works have studied the tidal deformation of NSs in binaries in the static limit, starting with Refs. [4, 5, 6]. The secular imprint on the GW signal associated to this tidal interaction is characterized by the star’s tidal deformability. The GW170817 event placed a constraint on the binary’s mass-weighted average tidal deformability, disfavoring extremely stiff EOSs [28, 7, 29, 30, 31]. Combined with lower bounds on the maximum mass from radio pulsar timing [32, 33, 34, 35] and joint mass–radius constraints from NICER X-ray timing [36, 37, 38, 39, 40, 41, 42, 43], these observations form the current picture of EOS constraints from the astrophysical side.
@@
The tidal response characterized by the tidal deformability is only the leading order effect obtained under the assumption that the tidal interaction is adiabatic. As the binary approaches the merger, this assumption gradually breaks down, and dynamical, frequency-dependent contributions in the tidal response function must be taken into account [13, 14, 44, 45]. In particular, NSs respond dynamically to the tidal field through the excitation of various oscillation modes as the orbital frequency sweeps through the detector band [10, 11, 12, 46, 47, 48, 49]. Therefore, dynamical tidal interactions provide a realization of NS asteroseismology [50, 51, 52]. That is, the study of the interior composition the star through its oscillation spectrum.
@@
The interior of NSs is more complicated than what can be described with just a prescribed density-pressure relation. Matter densities span by more than 14 orders of magnitude from the surface to the core, accompanied by changes in composition, from ordinary baryons to possibly exotic particles such as hyperons or quarks, as well as by distinct phase structures, such as solid and superfluid. Many low-frequency modes (compared to $f$-mode) restored by different physical mechanism may become tidally excited. Examples include gravity ($g$-) modes that arise from compositional gradients [10, 61, 62, 63, 64, 65, 66, 67], interface ($i$-) modes associated with discontinuities in density [68, 69, 70, 67, 71] or shear modulus [46, 72, 73, 71], and shear ($s$-) modes restored by shear stresses of the solid crust [74].
The tidal coupling of the foregoing modes is typically much smaller than that of the $f$-mode, but their lower frequencies make resonance more accessible earlier in the inspiral, and thus potentially detectable with current GW detectors. Several works showed that the GW phase shift induced by compositional $g$-modes is generally smaller than $10^{-3}$ radians; see, e.g., Refs. [62, 64, 75]. However, Refs. [76, 77] suggested that nonlinear effects could lock the $g$-mode resonance resulting in several radians of dephasing. Moreover, Refs. [46, 72, 73, 68, 69, 70, 73] suggested that the $i$-mode associated with the crust-core interface or a first-order phase transition may produce approximately $1$ to $10$ radians of GW dephasing. If observed, this dephasing would serve as a probe into the existence of phase transitions to quark matter, as well as into nuclear-matter parameters near the saturation density.
@@ II.1.1 Adiabatic index and stratification
$\quad\left.+\left.\frac{\partial\,(\mu_{e}-\mu_{\mu})}{\partial n_{\mathrm{b}}}\right|_{Y_{i}}\,\frac{\mathrm{d}Y_{\mu}}{\mathrm{d}n_{\mathrm{b}}}\right]\,.$ (4)
@@ IV.1 $i$-mode of nonstratified NSs
We show the $i$-mode’s eigenfunctions in the bottom panel of Fig. 6 for the canonical model. The radial displacement $\xi^{r}$ decays into both the core and the crust, as has a characteristic kink-like structure at the core-crust interface. The tangential displacement $\xi^{\theta}$ is significantly excited in the crust but is highly suppressed in the core due to the abrupt change in shear modulus, resulting in a discontinuity at the interface. Since the stress is related to the spatial derivative of displacements, the large displacement gradient at the the base of the crust can induce large shear stresses there. Some literature (e.g., Refs. [46, 72]) thus suggests that the resonantly excited $i$-mode can potentially yield the crust breaking if amplitude.
@@ V.1 Energy transfer and estimates of GWs
Figure 14: The relation between the mode frequency $f_{\alpha}$ and the tidal overlap integral $Q_{n2}$ for models with masses from $0.4\,M_{\odot}$ up to the maximum mass of $2.08\,M_{\odot}$. The circles and the squares denote the cases for $0.4\,M_{\odot}$ and $1.35\,M_{\odot}$, respectively. Also plotted are some representative contour lines of $E_{\rm mode}^{\rm max}$ [Section V.1] and $\Delta\Phi$ [Section V.1].
@@
Figure 16: The distribution of the strain $\sigma$ for different modes based on the canonical model. The upper panel presents the results for high-frequency modes—$f$, $s_{1}$, and $s_{2}$ at the time when $g_{1}$-mode reaches resonance. The lower panel displays the results for low-frequency modes, including the $g_{1}$ and $g_{2}$ modes, as well as the $i$-mode in the $\Gamma=\Gamma_{0}$ case. The black dashed lines indicate the location of the neutron drip. The companion is located at the direction of $\theta=90^{\circ}$.
@@ V.2 Crust breaking and the von Mises criterion
Remember again that $T_{1}$ and $T_{2}$ are the radial and tangential tractions, defined in Eq. (25); $V$ arises in the angular components of the displacement vector, defined in Eq. (12); $\lambda$ is a background metric function; and $\mu$ the shear modulus. We note that our expression for the strain differs from that in Ref. [141], Eq. (58), although the source of this discrepancy is not yet understood. For reference, we have made a Mathematica notebook with our derivation available online [162].
In Fig. 16, we illustrate the strain amplitude for different oscillation modes of the canonical model. Since the $f$-mode does not undergo resonance, and the $s_{1}$- and $s_{2}$-modes reach the resonance point very close to merger, we present the results for these high-frequency modes at the amplitude $|q^{\alpha}|$ corresponding to the orbital phase at which the $g_{1}$-mode is in resonance. In contrast, the results for the low-frequency modes ($g_{1}$-, $g_{2}$-, and $i$-modes) are shown at their respective resonance points, with amplitudes evaluated at resonance. Except the $i$-mode, a general feature is that the strain is largest along the direction to the companion ($\theta=90^{\circ}$) compared to other $\theta$ values. Owing to the linear nature of the problem, the strain amplitude at any other orbital phase can be obtained by rescaling with the value of $|q^{\alpha}|$ provided in Fig. 17.
@@
Here, the amplitude of the $f$-mode is set to its value at the $g_{1}$-mode resonance. In Fig. 18, we plot the radial profile of $rV^{\prime}$ within the crust for a $1.35\,M_{\odot}$ NS. Combined with the estimate in Section V.2, this analysis indicates that the outer crust is already broken in the vicinity of the $g_{1}$ resonance. Meanwhile, the peak-valley structure in the inner crust region observed in Fig. 16 arises from the radial variation of $V^{\prime}$, as illustrated in Fig. 18. It is worth noting that static tides are generally ineffective at fracturing the crust, as demonstrated in [165, 166]. This is primarily because the solid crust deforms in a manner similar to a fluid under static tidal forces.
Figure 18: Main contribution to the crustal strain ($r\,|{\rm d}V/{\rm d}r|$; see main text) as a function of $r$ for the $f$-mode of a $1.35\,M_{\odot}$ NS.
@@ VI Summary and discussions
Our work represents a step toward more realistic modeling of dynamical tides in binary NSs. In this study, we mainly establish the formalism using a single EOS, and there remains a lot to be done. For example, as is well known, the frequencies and tidal overlap integrals of $g$-, $i$-, and $s$-modes depend sensitively on nuclear parameters near saturation density. A systematic investigation of these dependencies is crucial for clarifying the disappearance of the $i$-mode, verifying the tensions in the predicted frequencies and tidal overlaps of $i$-modes compared to Refs. [46, 72, 73], assessing the role of $g$-mode resonances, and evaluating crust breaking triggered by both resonant and nonresonant excitations.
It is also important to note that further advances are needed in both the microphysical description of matter and the relativistic formulation of the problem. More specifically, for matter fields, superfluidity in the outer core is expected to increase the frequencies of the gravity modes [172, 171, 62], potentially modifying the mode spectrum and the crustal breaking behavior predicted in our work. Going beyond $npe\mu$ matter studied here, phase transitions to exotic degrees of freedom — such as hyperons or deconfined quarks — may lead to new phase structures, accompanied by additional $g$- and $i$-modes [173, 66, 68, 69, 174]. Recent studies suggest that $i$-mode resonances associated with first-order phase transitions can produce significant phase shifts in GWs, provided that such phase transitions do exist. In addition, bulk viscosity may be important for exotic phases of matter and could potentially result in appreciable tidal heating during resonances [175, 176, 177]. Regarding the modeling of dynamical tides, our treatment of the tidal coupling is based on a relativistic extension of the Newtonian formalism. It would be interesting to repeat our calculations in a full self-consistent relativistic approach (e.g., Refs. [14, 45, 178, 44, 179, 180, 181] for recent progress). Finally, nonlinear effects induced by tidal resonances may also lead to prolonged mode excitations, amplified GW phase shifts [76, 77, 152], and induce differential fluid flows within the NS interior [153, 60]. We will investigate these aspects in future work.
@@ Acknowledgements
Y. G. thanks Zexin Hu, Jan Steinhoff, and Yanbei Chen for useful discussions. H. O. S. acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) - Project No.: 386119226. Numerical computations were performed on the Sakura cluster at the Max Planck Computing and Data Facility. This work was in part supported by Grant-in-Aid for Scientific Research (grant No. 23H04900) of Japanese MEXT/JSPS.
@@ A.1.1 Perturbation equations for perfect fluid
${K^{\prime}}$ $=\frac{1}{r}H_{0}+\frac{n+1}{r}H_{1}+\left(\frac{1}{2}\nu^{\prime}-\frac{\ell+1}{r}\right)K-\frac{8\pi\,\left(\epsilon+p\right)}{r}e^{\lambda/2}W\,$, (63)
$W^{\prime}$ $=-\left(\frac{\ell+1}{r}+\frac{p^{\prime}}{\Gamma_{1}p}\right)W+re^{\lambda/2}\left[\frac{\epsilon+p}{\Gamma_{1}p}\left(e^{-\nu}\omega^{2}V+\frac{1}{2}H_{0}\right)-\frac{2(n+1)}{r^{2}}V+\frac{1}{2}H_{0}+K\right]\,$, (64)
$X^{\prime}=$ $-\frac{\ell}{r}X+\frac{1}{2}(\epsilon+p)e^{\nu/2}\left\{\left[\frac{1}{r}-\frac{\nu^{\prime}}{2}\right]H_{0}+\left[r\omega^{2}e^{-\nu}+\frac{n+1}{r}\right]H_{1}\right$.
$+\left[\frac{3}{2}\nu^{\prime}-\frac{1}{r}\right]K-\nu^{\prime}\frac{\ell(\ell+1)}{r}V-\frac{1}{r}\left[8\pi(\epsilon+p)e^{\lambda/2}+2\omega^{2}e^{\lambda/2-\nu}\right.\left.\left.-\,r^{2}\left(r^{-2}e^{-\lambda/2}\nu^{\prime}\right)^{\prime}\right]W\right\},$ (65)
@@
$\omega^{2}(\rho+p)V=e^{\nu/2}X+\frac{1}{r}p^{\prime}e^{\nu-\lambda/2}W-\frac{1}{2}(\rho+p)e^{\nu}H_{0}\,.$ (66)
@@
$V^{\prime}=\left(-A+\nu^{\prime}-\frac{\ell}{r}\right)V+\frac{Ae^{\nu}}{2\omega^{2}}\left(H_{0}+\frac{\nu^{\prime}}{re^{\lambda/2}}W\right)+rH_{1}-\frac{e^{\lambda/2}}{r}W\,.$ (69)
As a summary, in the fluid core, we use Eqs. 62, 63, 64, 65, 66, A.1.1 and 68 if the oscillation frequency $\omega\geq 0.01/M$. Otherwise, we use Eq. 69 to replace Eq. 65. For the fluid ocean (if added), the systematic equations Eqs. 62, 63, 64, 65, 66, A.1.1 and 68 are applied.

### v2

@@ I Introduction
As the densest stable form of matter known in the Universe, neutron stars (NSs) are expected to have complex interiors dominated by strong interactions inaccessible through terrestrial experiments. Despite extensive experimental and theoretical efforts in probing the nuclear equation of state (EOS), our current knowledge about the EOS at the densities most relevant for determining the internal structure of NSs remains poor [102, 122, 19].
@@
Compact binary mergers involving NSs provide multiple pathways to probe the dense matter, from subtle finite-size effects in the late inspiral [65, 43, 3, 63, 71, 135, 150, 101, 83, 64, 155] to the dramatic dynamics of merger remnants [149, 162, 18, 23, 38, 70], tidal disruption [99, 100, 98], and mass ejection [69, 81, 148]. The landmark detection of gravitational waves (GWs) from the binary NS inspiral GW170817 by the LIGO-Virgo-KAGRA Collaboration [3], along with its postmerger electromagnetic (EM) counterparts [4], marked the beginning of GW multimessenger astronomy, and has helped shed a new light on the EOS problem. However, the constraints on the EOS set by GW170817 are limited by the insufficient sensitivity of current GW detectors at the frequency band of the late-inspiral and merger, where matter effects on the GW waveform become most salient.
@@
An upgrade to the current ground-based detectors is pending for the coming years, and the next observing run of the International Gravitational-Wave Observatory Network, tentatively planned for 2028, will provide an opportunity to detect GW signals emanated during the merger from GW170817-like events. A central question to be addressed is then: To what extent can observations of binary mergers truly reveal the composition and state of matter inside NSs? Because a NS’s characteristic oscillation frequencies are determined by its internal structure, identifying quasiperiodic oscillations (QPOs) in some observations offers a way to probe these extreme environments. This article focuses on a subset of the stellar spectrum that may imprint observable signatures in gravitational waveforms or EM flares through tidal interactions.
@@
Much work has focused on the tidal deformation of NSs during binary interaction in the static limit [65, 43, 63]. The secular imprint on the GW signal associated to it is characterized by the so-called tidal deformability. The GW170817 event placed a constraint on the binary’s mass-weighted average tidal deformability, disfavoring extremely stiff EOSs [15, 3, 33, 185, 49]. Combined with lower bounds on the maximum mass from radio pulsar timing [34, 16, 44, 32] and joint mass–radius constraints from NICER X-ray timing [113, 112, 141], these observations form the current picture of EOS constraints from the astrophysical side.
@@
The tidal response characterized by the tidal deformability is only the leading order effect obtained under the assumption that the tidal interaction is adiabatic. As the binary approaches the merger, this assumption gradually breaks down, and dynamical, frequency-dependent contributions in the tidal response function must be taken into account [64, 155, 130, 62]. In particular, NSs respond dynamically to the tidal field through the excitation of various oscillation modes as the orbital frequency sweeps through the detector band [150, 101, 83, 170, 28, 94]. Therefore, dynamical tidal interactions provide a realization of NS asteroseismology [13, 14, 7]. That is, the study of the interior composition the star through its oscillation spectrum.
@@
The interior of NSs is more complicated than a prescribed density-pressure relation. Matter densities span more than 14 orders of magnitude from the surface to the core, accompanied by changes in composition, from ordinary baryons to possibly exotic particles such as hyperons or quarks, as well as by distinct phase structures, such as solid and superfluid. Many low-frequency modes (compared to $f$-mode) restored by different physical mechanism may become tidally excited. Examples include gravity ($g$-) modes that arise from compositional gradients [150, 140, 180, 104, 31, 183, 182, 181], interface ($i$-) modes associated with discontinuities in density [111, 30, 128, 181, 103] or shear modulus [170, 123, 186, 103], and shear ($s$-) modes restored by shear stresses of the solid crust [86].
The tidal coupling of the foregoing modes is typically much smaller than that of the $f$-mode, but their lower frequencies make resonance more accessible earlier in the inspiral, and thus potentially detectable with current GW detectors. Several works showed that the GW phase shift induced by compositional $g$-modes is generally smaller than $10^{-3}$ radians; see, e.g., Refs. [180, 31, 93]. However, Refs. [97, 96] suggested that nonlinear effects could lock the $g$-mode resonance resulting in several radians of dephasing. Moreover, Refs. [170, 123, 186, 111, 30, 128, 186] suggested that the $i$-mode associated with the crust-core interface or a first-order phase transition may produce approximately $1$ to $10$ radians of GW dephasing. This offers a promising probe of the existence of phase transitions to quark matter, as well as into nuclear parameters around the saturation density.
@@ II.1.1 Adiabatic index and stratification
$\left.+\left.\frac{\partial\,(\mu_{e}-\mu_{\mu})}{\partial n_{\mathrm{b}}}\right|_{Y_{i}}\,\frac{\mathrm{d}Y_{\mu}}{\mathrm{d}n_{\mathrm{b}}}\right]\,.$ (4)
@@ IV.1 $i$-mode of nonstratified NSs
We depict $i$-mode’s eigenfunctions in the bottom panel of Fig. 6 for the canonical model. The radial displacement $\xi^{r}$ decays into both the core and the crust, as has a characteristic kink-like structure at the core-crust interface. The tangential displacement $\xi^{\theta}$ is significantly excited in the crust but is highly suppressed in the core due to the abrupt change in shear modulus, resulting in a discontinuity at the interface. Since the stress is related to the spatial derivative of displacements, the large displacement gradient at the the base of the crust can induce large shear stresses there. Some literature (e.g., Refs. [170, 123]) thus suggests that the resonantly excited $i$-mode can potentially yield the crust breaking if the amplitude is large enough.
@@ V.1 Energy transfer and estimates of GWs
Figure 14: The relation between the mode frequency $f_{\alpha}$ and the tidal overlap integral $Q_{n2}$ for models with masses from $0.4\,M_{\odot}$ up to the maximum mass of $2.08\,M_{\odot}$. The circles and the stars denote the cases for $0.4\,M_{\odot}$ and $1.35\,M_{\odot}$, respectively. Also plotted are some representative contour lines of $E_{\rm mode}^{\rm max}$ [Section V.1] and $\Delta\Phi$ [Section V.1].
@@
Figure 16: The distribution of the strain $\sigma$ for different modes based on the canonical model. The upper panel presents the results for high-frequency modes—$f$, $s_{1}$, and $s_{2}$ at the time when $g_{1}$-mode reaches resonance. The lower panel displays the results for low-frequency modes, including the $g_{1}$ and $g_{2}$ modes, as well as the $i$-mode in the $\Gamma=\Gamma_{0}$ case. The black dashed lines indicate the location of the neutron drip. The companion is located at the direction of $\theta=90^{\circ}$. An animation of the strain evolution as a function of the inspiral phase can be found at [1].
@@ V.2 Crust breaking and the von Mises criterion
Remember again that $T_{1}$ and $T_{2}$ are the radial and tangential tractions, defined in Eq. (25); $V$ arises in the angular components of the displacement vector, defined in Eq. (12); $\lambda$ is a background metric function; and $\mu$ the shear modulus. We note that our expression for the strain differs from that in Ref.[124] (Eq.(58)), although the source of this discrepancy is not yet understood. For reference, we have made a Mathematica notebook with our derivation available online [2].
In Fig. 16, we illustrate the strain amplitude for different oscillation modes of the canonical model. Since the $f$-mode does not undergo resonance, and the $s_{1}$- and $s_{2}$-modes reach the resonance point very close to merger, we present the results for these high-frequency modes at the amplitude $|q^{\alpha}|$ corresponding to the orbital phase at which the $g_{1}$-mode is in resonance. In contrast, the results for the low-frequency modes ($g_{1}$-, $g_{2}$-, and $i$-modes) are shown at their respective resonance points, with amplitudes evaluated at resonance. Except the $i$-mode, a general feature is that the strain is largest along the direction to the companion ($\theta=90^{\circ}$) compared to other $\theta$ values. Owing to the linear nature of the problem, the strain amplitude at any other orbital phase can be obtained by rescaling with the value of $|q^{\alpha}|$ provided in Fig. 17. The time before merger is related to the frequency of GW as
$t-t_{\rm merger}\simeq-2.29\mathrm{~s}\left(\frac{1.175\,M_{\odot}}{M_{c}}\right)^{5/3}\left(\frac{100\,\mathrm{Hz}}{f_{\mathrm{gw}}}\right)^{8/3}\,.$ (59)
Here $t_{\rm merger}$ denotes the merger time, and $M_{c}$ is the chirp mass of the binary, which equals to $1.175\,M_{\odot}$ for an equal-mass binary NS with $M^{\prime}=M=1.35\,M_{\odot}$.
@@
Here, the amplitude of the $f$-mode is set to its value at the $g_{1}$-mode resonance. In Fig. 18, we plot the radial profile of $rV^{\prime}$ within the crust for a $1.35\,M_{\odot}$ NS. Combined with the estimate in Section V.2, this analysis indicates that the outer crust is already broken in the vicinity of the $g_{1}$ resonance. Meanwhile, the peak-valley structure in the inner crust region observed in Fig. 16 arises from the radial variation of $V^{\prime}$, as illustrated in Fig. 18. It is worth noting that static tides are generally ineffective at fracturing the crust, as demonstrated in [127, 55]. This is primarily because, under static tidal forces, the solid crust deforms much like a fluid. As a result, the star can experience significant global deformation while the internal shear strain remains small, making static tides inefficient at breaking the crust.
Figure 18: Main contribution to the crustal strain ($r\,{\rm d}V/{\rm d}r$; see main text) as a function of $r$ for the $f$-mode of a $1.35\,M_{\odot}$ NS.
@@ VI Summary and discussions
Our work represents a step toward more realistic modeling of dynamical tides in binary NSs. In this study, we mainly establish the formalism using a single EOS, and there remains a lot to be done. For example, as is well known, the frequencies and tidal overlap integrals of $g$-, $i$-, and $s$-modes depend sensitively on nuclear parameters near saturation density. A systematic investigation of these dependencies is crucial for clarifying the disappearance of the $i$-mode, verifying the tensions in the predicted frequencies and tidal overlaps of $i$-modes compared to Refs. [170, 123, 186], assessing the role of $g$-mode resonances, and evaluating crust breaking triggered by both resonant and nonresonant excitations. It is worthwhile to note that, once the crustal strain exceeds the elastic limit, plastic flow may occur, converting oscillation energy into thermal one [20, 82, 123, 186]. It will be interesting to examine how different oscillation modes contribute to such heating.
It is also important to note that further advances are needed in both the microphysical description of matter and the relativistic formulation of the problem. More specifically, for matter fields, superfluidity in the outer core is expected to increase the frequencies of the gravity modes [78, 125, 180], potentially modifying the mode spectrum and the crustal breaking behavior predicted in our work. Going beyond $npe\mu$ matter studied here, phase transitions to exotic degrees of freedom — such as hyperons or deconfined quarks — may lead to new phase structures, accompanied by additional $g$- and $i$-modes [167, 182, 111, 30, 179]. Recent studies suggest that $i$-mode resonances associated with first-order phase transitions can produce significant phase shifts in GWs, provided that such phase transitions do exist. In addition, bulk viscosity may be important for exotic phases of matter and could potentially result in appreciable tidal heating during resonances [52, 53, 142]. Regarding the modeling of dynamical tides, our treatment of the tidal coupling is based on a relativistic extension of the Newtonian formalism. It would be interesting to repeat our calculations in a full self-consistent relativistic approach (e.g., Refs. [155, 62, 61, 130, 110, 60] for recent progress). Finally, nonlinear effects induced by tidal resonances may also lead to prolonged mode excitations, amplified GW phase shifts [97, 96, 131], and induce differential fluid flows within the NS interior [137, 90]. We will investigate these aspects in future work.
@@ Acknowledgements
We thank the anonymous referee for comments and suggestions. Y. G. thanks Zexin Hu, Jan Steinhoff, and Yanbei Chen for useful discussions. H. O. S. acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) - Project No.: 386119226. Numerical computations were performed on the Sakura cluster at the Max Planck Computing and Data Facility. This work was in part supported by Grant-in-Aid for Scientific Research (grant No. 23H04900) of Japanese MEXT/JSPS.
@@ A.1.1 Perturbation equations for perfect fluid
${K^{\prime}}$ $=\frac{1}{r}H_{0}+\frac{n_{\ell}+1}{r}H_{1}+\left(\frac{1}{2}\nu^{\prime}-\frac{\ell+1}{r}\right)K-\frac{8\pi\,\left(\epsilon+p\right)}{r}e^{\lambda/2}W\,$, (64)
$W^{\prime}$ $=-\left(\frac{\ell+1}{r}+\frac{p^{\prime}}{\Gamma_{1}p}\right)W+re^{\lambda/2}\left[\frac{\epsilon+p}{\Gamma_{1}p}\left(e^{-\nu}\omega^{2}V+\frac{1}{2}H_{0}\right)-\frac{2(n_{\ell}+1)}{r^{2}}V+\frac{1}{2}H_{0}+K\right]\,$, (65)
$X^{\prime}=$ $-\frac{\ell}{r}X+\frac{1}{2}(\epsilon+p)e^{\nu/2}\left\{\left[\frac{1}{r}-\frac{\nu^{\prime}}{2}\right]H_{0}+\left[r\omega^{2}e^{-\nu}+\frac{n_{\ell}+1}{r}\right]H_{1}\right$.
$+\left[\frac{3}{2}\nu^{\prime}-\frac{1}{r}\right]K-\nu^{\prime}\frac{\ell(\ell+1)}{r^{2}}V-\frac{1}{r}\left[8\pi(\epsilon+p)e^{\lambda/2}+2\omega^{2}e^{\lambda/2-\nu}\right.\left.\left.-\,r^{2}\left(r^{-2}e^{-\lambda/2}\nu^{\prime}\right)^{\prime}\right]W\right\},$ (66)
@@
$\omega^{2}(\epsilon+p)V=e^{\nu/2}X+\frac{1}{r}p^{\prime}e^{\nu-\lambda/2}W-\frac{1}{2}(\epsilon+p)e^{\nu}H_{0}\,.$ (67)
@@
$V^{\prime}=\left(-A+\nu^{\prime}-\frac{\ell}{r}\right)V-\frac{Ae^{\nu}}{2\omega^{2}}\left(H_{0}+\frac{\nu^{\prime}}{re^{\lambda/2}}W\right)+rH_{1}-\frac{e^{\lambda/2}}{r}W\,.$ (70)
As a summary, in the fluid core, we use Eqs. 63, 64, 65, 66, 67, A.1.1 and 69 if the oscillation frequency $\omega\geq 0.01/M$. Otherwise, we use Eq. 70 to replace Eq. 66. For the fluid ocean (if added), the systematic equations Eqs. 63, 64, 65, 66, 67, A.1.1 and 69 are applied.